OpenStudy (anonymous):

Find a polar equation for the curve represented by the given Cartesian equation. x^2-y^2=1

OpenStudy (anonymous):

Use: x = rcosθ and y = rsinθ. Then: x² + y² = 1 ==> (rcosθ)² + (rsinθ)² = 1 ==> r²cos²θ + r²sin²θ = 1 ==> r² * (cos²θ + sin²θ) = 1 ==> r² = 1 Then since r > 0, we can take the positive square root to get r = 1

OpenStudy (anonymous):

err wrong signs

OpenStudy (anonymous):

what do you mean wrong signs?

OpenStudy (anonymous):

u put -y^2 i accidentally put +

OpenStudy (anonymous):

how did you get ==> r² * (cos²θ + sin²θ) = 1 ==> r² = 1.. Do I just divide by the (cos²θ + sin²θ) to get 1? (like if i replace the plus sign with a subtraction sign)